3. Approximate the following integral using the two-point Gaussian quadrature rule 2 (x + a)?e(x-1)2-Bdx
Answer #1

Given. fon) S cata)? &CR-D2-3 In By Two point Gaussian quad ratione sule: 54cm) dn = Cs PC*) + Caf Cre) b where, a = b c Ca-*CX) = 460-4326) => (0:4226+w)? e60:4226-152 -33 (0.4226+)? f(x) = f ( J. 5713) (1.5773 tx)² e 2 (0.3333-3) (1.5793-182 - 3 @